True or false:


The value of the binomial coefficient "\displaystyle{{2}\choose{10}}" is zero.


1
Expert's answer
2021-09-22T01:01:40-0400

Solution:

We know that  "\displaystyle{{n}\choose{r}}=\dfrac{n!}{r!(n-r)!}" , where "n\ge r;n,r\ge 0"

But in  "\displaystyle{{2}\choose{10}}", "n=2,r=10"  which is against the definition of binomial coefficient.

Thus, the value of  "\displaystyle{{2}\choose{10}}" does not exist, neither it is zero.

Hence, given statement is false.